Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Sarah
Log entry time 17:53:58 on September 13, 2009
Entry number 289370
keyword=Run Coordinator's Report for September 12 Swing to September 13 Day
Run Coordinator's Report for September 12 Swing to September 13 Day
Summary
Not a good data-taking day.
Significant News
- The swing shift took data last night while the MCC and the
Compton experts tried to tune the beam for the Compton (to reduce the
background rates). They then did some harp scans to determine the beam
size, then data-taking resumed.
- The insertable half wave plate was put IN last night at 19:30, so we
are now on slug 6. However, directly after that the beam was lost in the
injector (see halog
entry 289289 and halog
289313 ). After a long night for the accelerator people, we got beam
back around 6:30am. We did diagnostics, lost the beam again, and got it
back. Production running began again about 10:00am.
- The Compton is still having issues, but the experts are working on
it.
Plans
- Take production data! Watch the lumi widths to see if anything
changes. Please check the histograms on the online data monitor with the
reference ones on the shift how-to page (http://hallaweb.jlab.org/experiment/HAPPEX/opmanual/opmanual.html#HPXStartingonlineanalysis).
- Take Compton data. Alexandre will start looking at the electron
detector signals. People will also take a look at the Compton data from
this morning, where we had both photon and electron data. They will not
be taking an access for the power supply.
- Run spot++ every couple of hours and log it. Change the prescale
factor PS8 to 1 and PS1 and 2 to 99999.
Concerns
- The Compton polarimeter.
- There are lots of empty shifts that need to be filled immediately!
Please sign up!
- Keep a close eye on the lumi widths as we take data.
- Please make sure that the shift checklist is done once each shift.
- The Flipper script didn't work the last time we did a half-wave
plate change, and so things had to be done by hand. This needs to be
fixed.
- A WAC schedule needs to be created and posted so that I know who is
doing this.
- Beam Time Accounting: take off 1 minute per beam trip, since the BTA
code doesn't.
- The dithering looks non-invertible. Someone should look at this.
- There is a constant alarm due to unused VDCs. This alarm should be
removed so that people don't become desensitized to the alarm handler.
Shift Summaries and BTA Reports